1. A hydrochloric acid solution is standardized using 0.500 g of sodium carbonate, Naze Find the molarity of the acid if 29.5 please correct of needed. thanks!
Answer #1

The balanced equation is

2 HCl(aq) + Na2CO3(s) ------------- 2 NaCl(aq) + CO2(g) + H2O(l)

2 mole         1 mole

mass of Na2CO3 = 0.500 g

molar mass of Na2CO3 = 105.99 g/mol

number of moles of Na2CO3 = mass of Na2CO3/molar mass = 0.500/105.99= 0.00472 moles

number of moles of Na2CO3 = 0.00472 moles

according to equation

1 mole of Na2CO3 = 2 moles of HCl

0.00472 moles of Na2CO3 = ?

                                          = 0.00472 x 2 /1 = 0.00944 moles of HCl

number of moles of HCl required = 0.00944 moles

volume of HCl = 29.50 mL = 0.02950L

molarity of HCl = number of moles of HCl / volume in L = 0.00944 / 0.02950 = 0.32 mol/L

molarity of HCl = 0.32M